OpenStudy (anonymous):

i think i figured this out correctly but not sure where to put the negative sign |-4/3|-|-5|/2 the answer i got is -7/6 is this correct and if so do i put the negative sign next to the seven cause when its lined up the negative sign is next to the division line hope that makes sence

myininaya (myininaya):

4/3-5/2 (8-15)/6 -7/6 yes thats right

OpenStudy (anonymous):

okay so on my answer sheet it has one box to input the numertaor and one box to input the denominator but there is nowhere to put the negative sign so what im not sure of is if the negative sign goes in the box with the7 or in the box with the 6 because when its written out it is 7 over 6 but the bar for over is where the negative sign is in the answer its lined up with that bar nbot with either number i hope that makes sence i dont know how else to explain it

myininaya (myininaya):

what? \[-\frac{7}{6}=\frac{-7}{6}=\frac{7}{-6}\] there are all the same but i'm not really sure what you are asking? maybe you are asking this i don't really know

OpenStudy (anonymous):

yes idk how u typed it like that but the 1st set of -7 over 6 see how the neagtive sign is lined with the diviion bar does that matter if its lined up like that

OpenStudy (anonymous):

The first way is the "prettiest" way to do it but any of them are satisfactory and they all mean the same.

OpenStudy (anonymous):

ok thanks for clarifying that for me i wasnt sure if where the negative sign was placed affected the answer